OpenStudy (anonymous):

1.What is the weight (in grams) of a liquid that exactly fills a 182.8 milliliter container if the density of the liquid is 0.135grams over milliliter? 2. A piece of aluminum occupies a volume of 13.0 milliliters and weighs 52.9 grams. What is its density of the aluminum rounded to the nearest hundredth?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

hint: Density = (Mass)/(Volume) or D = M/V

jimthompson5910 (jim_thompson5910):

# 1 D = M/V 0.135 = M/182.8 solve for M

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@sarah1017

OpenStudy (anonymous):

Some one please i only have these two i really don't know what to do that's why i asked.

OpenStudy (anonymous):

2. 4.0692307692308

OpenStudy (anonymous):

1.24678